NEED HELP WITH QUESTIONS E TO I1. Consumption :(-700 +0.95% Investment: I =500-30i Govt spending: (1=50 Money demandi L(i,Y)=0.75%-30i Money supply : M5/p=d) when G=100 , in new goods market equilibric y=700+ 0.gsyr500- 30irroo Y= 1200+0.95Y - 30i 0.05 Y = (300-301 9226000-600 i

Consider a hypothetical economy characterized by the following equations(all variables as defined in class).

Consumption: C = 700 + 0.95Y Investment: I=500− 30i
Government spending: G=50
Money demand: L(i,Y )=0.75Y − 30i Money supply: Ms/P=400

(a) What is the equation of the IS curve?

(b) What is the equation for the LM curve?

(c) Solve for the equilibrium values of income (Y) and interest rates (i).

(d) Assume that the government engages in expansionary fiscal policy by increasing

expenditure to G = 100. Solve for the new equilibrium values of income (Y) and interest

rates (i). Illustrate your answer in a graph.

(e) Assume that instead of the expansionary fiscal policy in part (d) above, that the Monetary

Authority (The Central Bank) engages in expansionary monetary policy by increasing the money supply by 100 to Ms/P = 500. Solve for the new equilibrium values of income (Y) and interest rates (i) in this case. Illustrate your answer in a graph.

(f) Whichofthepoliciesfromparts(d)and(e)abovewouldbethemosteffective expansionary policy? Explain.

(g) Say that the government engaged in both of the policies of parts (d) and (e) simultaneously. Would this policy mix be more or less expansionary than the individual policies?

(h) Suppose that money demand in this economy was observed to be income insensitive so that it was now represented by L(i, Y ) = L(i) = 600 − 30i. In this case, which of the policies from parts (d) and (e) above would be the most effective expansionary policy? Explain your answer in terms of the concept of “crowding out”.

(i) Lastly,returningtotheoriginalmodelasinpart(a),supposethatinvestmentwasnow determined to be interest insensitive such that it could be represented by I = 600. In this case, which of the policies from parts (d) and (e) above would be the most effective expansionary policy? Explain this answer in terms of “crowding out”.
Answer #1

gorernmentongay in pantiaay new Ms SOO P monekoy peley New LM cwme eguahon wil be- Liy) 0.75Y-801 -S00 S00 +30 Cso0 +30i 666.666-66 t1520.8 Neiw Y= 2/87.96 Graphially LM 38.22 38.02 IS 2062.8 2187Y6 The menehaoy poliy is mre gfecthe expansionary peliNew Is frem port dY 26000-600 Nw LM fom pat e) y 66-66+40 Equactrng tolth, ue get 666.66+907 26000-6001 26000-666.66 2S7333.3if mony Indensifve 20y) li 600-30 domand i interet The old Cwithood eanionary poley LM eguaton wil be 600-30 600 -4.00 30 6.6fer expantienacy monhy paly LM eguilibium naio 600-301 301 100 1= 30 ie 3.33 Nee , Coth old s nes LM/antexe 2 25G000-6001 2s,IS ALO Sold LMold S-33 21004 У 22,004 23.002 21,002findercest Ini ensthe meney In care domand M wne esa! horizental. with an expartionary iscal pody only GBP ,y will change asI case of insenthre S wrre ue a rextical S cune have ce eect b pobciel In the Thus, bue can chagram 1S neo neu E E3 Ez y eguiffken expantionary Ascal poly IS shif hoSnew and heuo eilbum s EL wrt espartionay monetary poty avd quilbrun shifts frem E 7గ